Circular No. 7491

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dj IN NGC 735 S. Benetti, A. Zacchei, I. Perez, M. Pedani, and A. Buzzoni, Telescopio

Nazionale Galileo (TNG), communicate: “A fully reduced CCD spectrum (range 350.4–812.7 nm, resolution 1.1 nm), obtained on Sept. 15.10 UT with the TNG reflector (+ D.o.lo.res spectrograph), shows SN 2000dj (cf. IAUC 7490) to be a type-II supernova around maximum light. The spectrum consists of a blue continuum with superimposed P-Cyg lines of Ca II H and K and hydrogen. The average expansion velocity deduced from the minima of the Balmer lines is 9800 km/s (a recession velocity for the parent galaxy of 4629 km/s has been obtained from NED database). The absorption of Hα line has a double profile with a secondary minimum measured at 629.1 nm.”

M. Armstrong, Rolvenden, Kent, England, reports his independent discovery of SN 2000dj at mag 17.0 on a 45-s CCD exposure taken with a 0.30-m reflector on Sept. 11.011 UT in the course of his supernova patrol (limiting mag 19.0). He gives the following position end figures (cf. IAUC 7490): 40^{s}.63, 12′′.9. The new object does not appear on images taken by Armstrong on 1999 Nov. 6 (limiting mag 19.5) or 2000 Aug. 28.987 (limiting mag 18.5), nor does it appear on Palomar Sky Survey images taken on 1992 Oct. 31 (limiting mag 22.5) and 1989 Nov. 29 (limiting mag 20.8)

V4633 SAGITTARII R. J. Rudy, D. K. Lynch, S. Mazuk, and C. Venturini, The Aerospace

Corporation; R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego; and T. Armstrong, Thomas K. Armstrong Development Corporation, report 0.8–2.5-μm spectrophotometry of V4633 Sgr = N Sgr 1998 using the Shane 3-m telescope and the Aerospace Near Infrared Imaging Spectrograph on July 21.36 UT, ∼ 850 days after peak brightness: “V4633 Sgr has faded by more than a magnitude in the past 11 months (see IAUC 7259). J, H, and K magnitudes determined from the spectrophotometry are 14.8, 14.4, and 13.8. He II at 1.0124 μm exceeds He I 1.0830-μm in strength, indicating that most of the helium is doubly ionized. The strongest features in the spectrum are the coronal lines of [Si VI] at 1.9629 μm and [Ca VIII] at 2.3214 μm. [Si VII] 2.4827-μm, and [S VIII] 0.9913-μm have appeared in the last 11 months. Surprisingly, the spectrum does not exhibit the large degree of reddening expected given the object’s galactic coordinates (l = 5^{o}.1, b = −6^{o}.2).”

2000 September 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green